OpenStudy (anonymous):

The last time you bought pizza, 3 pizzas were just enough for 7 people. At this rate, how many pizzas should you buy for a party of 33 people? a.7 pizzas b.11 pizzas c.13 pizzas d.15 pizzas

OpenStudy (lilg132):

7/3 = 2.333333 people per pizza 33 / (7/3) = 99/7 pizzas round it off 15 pizzas d. 15 pizzas

OpenStudy (anonymous):

medal thanks